OpenStudy (anonymous):

Jeff and Sharon drove for 4 hours with a wind speed of 10 mph. Traveling against the same wind speed, it took them 6 hours to make the return trip. How fast were they traveling in still air? Answer 60 mph 65 mph 55 mph 50 mph

OpenStudy (anonymous):

If x is her speed without wind, then x+10 is her speed with the wind, and x-10 is her speed against the wind. Using rate x time = distance, and since the distances are the same in both directions, we have 4(x+10)=6(x-10). Solving for x give 50 mph.

jimthompson5910 (jim_thompson5910):

Trip there: d=rt d=(x+10)*4 d=4x+40 Trip back: d=rt d=(x-10)*6 d=6x-60 -------------------- d=4x+40 6x-60=4x+40 6x=4x+40+60 6x-4x=40+60 2x=40+60 2x=100 x=(100)/(2 x=50 So their speed in still air is 50 mph